The number a is real and such that the equation x^2 +2 (a - 1) x - a + 7 = 0 has

  two different real negative solutions. One can then conclude that

  (a) a <−2; (b) 3 <a <7; (c) it is impossible; (d) none of (a) - (c).


1
Expert's answer
2021-05-07T14:23:37-0400

Given equation-

"x^2+2(a-1)x-a+7=0"


Let the required roots be -m and -n


then sum of roots "-m-n=-2(a-1)~~~~-(1)"


Also Product of roots "-m\times -n=-a+7~~~~~-(2)"


"m+n=2a-2 \text{ and } mn=7-a"


As, "mn>0\Rightarrow 7-a>0\Rightarrow a<7"


Also, "m+n>0\Rightarrow 2a-2>0\Rightarrow a>1"


So "1<a<7."